Mr.Thompson invests a principal of $8000; part at 8% per annum and the rest at 10% per annum. After one year, the total interest was $724. How much did Mr. Thompson invest at each rate?
Answer by checkley71(8403)

.10X+.08(8000-X)=724
.10X+640-.08X=724
.02X=724-640
.02X=84
X=84/.02
X=4200 INVESTED @ 10%
8000-4200=3800 INVESTED @ 8%
PROOF
4200*.10+3800*.08=724
420+304=724
724=724